OpenStudy (anonymous):

Find an equation of variation where y varies directly as x and y = 10 when x = 5, find y when x is 4.

OpenStudy (akashdeepdeb):

If y increases x increases that's what this means so! y1/x1 = y2/x2 [Another way of writing proportionalities varying directly] SO 10/5 = y2/4 y = 8 Understood? :)

OpenStudy (anonymous):

yessssssssss